Back to the NIKA2 run 6 main page
October 21, (Friday)
Martino arrives at the telescope in the evening. Alessandro, in remote, has been optimizing the way to put some pressure in the mixture circuit in order to reach the (local) atmospheric pressure, ie about 720mbar. The solution he proposes, switching of one of the two PT compressor, works very well. Bravo! We also have to get a bit of mixture back in the reserve, roughly 70mbar, otherwise the pressure increases above the desired value. Martino mounts the accelerometer on the NIKA2 'nose', but the reading doesn't seem sensitive enough to show clearly the PT induced vibrations -> solution for the moment not implemented! In the meantime, Alessandro restarts the second compressor to leave the system in precool mode overnight. So tomorrow the cryostat will be nicely cold at ~5K.
October 22, (Saturday)
Yesterday evening at 23 the telescope has been stopped due to the wind. This leaves us more freedom to play in the cabin. So we (= Martino in the cabin and Alessandro remotely connected, taking care of the cryostat and sharing the responsability..) begin the realignement process, the basically implies:
- Increase the pressure in the precool circuit to be able to move the PT head. This is achieved switching off one PT compressor.
- Move the head (one at a time), feel where it starts to touch to the NIKA2 4K, and find the ideal 'central position'.
- Holding it in this position, screw the dumping screws until the touch and tighten them a little.
- Restart the PT compressor, go back in the precool phase and when it's time launch the condensation
- During condensation, optimize the PT2 position by playing (gently!) with the screws and looking for a max in the 4K plate temperature. NOTE: typically starting from the ideal position, turning the screw by 45° in any direction is enough to start touching. So it's a very delicate process!!
- When the condensation is finished and the system is cold, check on the Camadia time traces that the noise is ok.
To give an idea of timing, today we did:
7h45 Switch off 1 PT compressor
8h40 P in the 'gaines' is ok, start to play around with PT head position
9h45 PT position adjusted, restart the standard cooldown by switching back on the second PT compressor.
12h00 Launch condensation. NOTE that we could do this very soon because the system had been already well precooled before starting the whole PT adjustement process. If it's not the case, wait for the correct temperatures beafore launching the condensation!
20h00 Tbm at ~200mK, good enough to see how things are going and if the PT lines have disappeared (but not ok for detailed tests as T is not yet stable enough)
October 23, (Sunday)
The noise traces seem to indicate that the issue has been solved. Playing around with the tone powers and with camadia to get a feeling of its current configuration. It seems to be pretty stable, very well! Still, too soon to give definitive answer: we'll have to see what goes on when the telescope moves (it's still stopped do to the wind, constantly very high!), and especially when we go on sky.
October 24, (Monday)
Still windy! It's now more then 48hrs of continuous 'wind stop'. Tonight the gusts reached 200km/h.. This morning wind, rain and hail. The gusts were still too strong and the transport has been canceled: Juan M-P, that was coming up with it, had to go back to Granada and will arrive tomorrow with the rest of the NIKA2 crew. Hopefully the weather will improve. It should, but the tuesday to wednesday night might still be affected by the wind. Let's hope not..
Created a 'run19.ini' parameter: it should in theory be identical to run18.ini, apart from the power on each tone, slightly changed (in general to lower values). The noise spectra seem correct but better to check on the sky. Plus, sometimes there seems to appear some noise (similar to PT) but not constantly. Can it be related to the cryostat? I have noticed that the whole cabin is (slightly) saking when the wind gusts are strong, and as a consequence, the whole NIKA2 cryostat start to move, in a quite impressive way.. So better to check the PT effect when there is a calmer atmosphere. Especially considering that the noise spectra are already quite good. Better not to risk worsening the things if the source of the problem is not completely sure that it is PT related..
October 25, (Tuesday)
Xavier, Florian, Helene, Charles and Juan arrived. Very windy. Many Internet problems during maintenance. We only did few simulated observations to test new elvin code procedures.
The shifts were divided into: * Juan and Florian: start of observations until 2am * Xavier, Helene, and Charles: 2am until 8am
October 26, (Wednesday)
Started observations in particular poor weather conditions and still a bit windy - some dark tests. TAPAS is not working.
In the afternoon, we start observing at 18h as expected. Wind has calmed down but opacity is really high (>0.7). The sky is actually warmer than the window! Anyway, we can start to make pointings, focuses, and beam maps. The atmosphere is pretty stable (but degrading a bit over time). The system seems to be working pretty well, and no PT lines are observed.
NOTE: at around 23h local time, we decreas the power on the first band of the four lines of array2. They looked a bit over-driven, especially for feed4 that had problems tuning correctly the KIDs. So we lower them, updating the run19.ini file, from 30-30-30-30 to 25-25-25-20 (ordered AMC A-B-C-D).
October 27, (Thursday)
The second shift of the night went smoothly (although still in the company of a very thick atmosphere..). But the system is pretty stable.
Robert arrived at the telescope today.
In the morning, a moment of panic when a clear ray of noise in the Camadia noise spectra appears at 1.4Hz. Fearing that it might be a sign of the PT starting to touch, Martino plays a bit with the screws of the PT head -> no big changes on noise, a slight change in the 4Kplate temperature, that increases a bit (-> in theory, a good thing!). Then, we notice that there is a big pickup when the chopper of Emir is turning: the window is open and looking at 300K! So we put back the alu flange to close the window --> the 1.4Hz noise line disappears! So, BE CAREFUL with this line! If you see it, before starting to despair, check what is the status of the window and, if open, close it and try again! Still, it would be nice to investigate the origin of such a pickup line.. Anyway, now with the window closed the noise traces are similar to those found yesterday, so it should be fine!
Martino leaves at around 16h, hoping that NIKA2 will keep behaving as it has done until now during the week.. Do not hesitate to call any of us in case of need, the numbers are on a sticker on the mac8 screen. Good luck... and let's hope that the tau will finally decrease!
October 28, (Friday)
Given that Martino left, the Turbo Pump encountered issues. Martino was called, and the cooling rate was reduced to 30%. This left the array at 205 mK. The resolution took some time, but the weather was still bad (primarily in terms of opacity). Observations started with dark scans. Once on sky, focus maps were performed, and then beammaps on Uranus and, in the second shift, beammaps on 3c84 were performed. We also verified that 2251+158 is about 3 Jy in 1mm (as opposed to 3c84 at ~8 Jy) and verified the ability to track and point to small solar system objects (i.e. asteroids) by observing Vesta.
Early analysis Friday morning/afternoon suggests that the higher array temperature translates into slightly noisier pixels.
October 29, (Saturday)
The weather started off poor Friday afternoon/evening, but quite rapidly improved (tau ~ 0.2 by the end of the night). This made early observations difficult as conditions were rapidly changing (but for the better). Observations went well, mainly in focus and defocus beammaps on Uranus ...
TAPAS data filler had to be restarted and started to work towards the start of this shift. However, comments could not be entered.
Just before the transition between the shifts, a focus in all directions was done on Uranus, and after the shift, several beammaps on Uranus were performed, and a quick jaunt to Ceres. We then look to observe several secondary calibrators, but want to first stop at 3c84 and refocus. Around 3:00 UTC, the acquisition stopped working. In practice, this was noticed when trying to reduce the first of the focus scans (20161029s21), and IDL could not find the data file. The problem was soon found (at least, on the surface) to be a lack of connection with Elvin (it was not flashing in Acqui). We tried a relaunch of the Acqui and related windows, and we saw Elvin flashing, but did not seem to get the rest of the system back to working. After a couple more tries, and with the assistance of Juan, the solution was to do a hard reboot of the crate of array 2, and upon relaunching Acqui, everything appeared to work. We note here, that just before the focus scans, the telescope operator, Ignacio, had asked us to not submit any scans, and he followed his given instructions/protocol to restart certain systems to get the telescope alarm webpage working again properly. Among the systems restarted was Elvin, which appeared to have restarted properly as reported in the terminal where he had typed his commands. However, there is no reason why the issue with crate 2 should be related to Elvin. By 4:00 UTC, Acqui appeared to have connection with all other computers/boards. However, there were some difficulties retuning the detectors. We try to start where we were before just before 4:30 UTC, but we are missing several resonances in array 2, and the array is noisy. Another reboot (soft, or cold, reboot this time, from nika2-a) of the crate was performed, but no improvement in the resonances were found. Changing the input power to the detectors also does not seem to improve the results. With ~1 hour left of potential observations, but a non-optimal array 2, we ceded the remainder of the time to the next project. However, in the time spent attempting to contact the next project (which was, in the end, impossible), another reboot of the crate (2) "magically" resolved the issue, and we were able to regain most resonances and the noise in the detectors were as they should be. We continued just with a pointing, Z-focus, and beammap99sub. However, during the beammap, it appears as though we are not that close to the detector resonances.
October 30, (Sunday)
The first half of the night was primarily focused on observing secondary calibrators; all of the ones on the "to-do" list were done. Observations then shifted to out of focus beammaps. At the transition of shifts, we were nearing completion of the series with 5 different X focii on Uranus. In the second shift, we finished the series of X focii and did a series of 5 different Y focii. After beammaps, we moved to 3C84 for pointing and focus, where we (again) experienced difficulties; this time limited only to the telescope, but it still cost us about an hour. We proceeded to NGC 891 for mapping / testing consistency. Then on to other secondary calibrators and a skydip, including a test of a faint source: Lutetia.
October 31, (Monday)
At 22h50 UT (30-10-2016) we had a problem with the acquisition. Everything seemed to be working well (the arrows were well aligned, the electronic boxes tag were blinking and the histograms were fine) but there were actually error messages (Qt errors) in a terminal. We therefore lost 2 scans (one corrupted but with the acquisition working and one stopped because the acquisition crashed). We therefore rebooted the acquisition but there was a problem with the synthesizer of array 2. We therefore went up in the cabin to reboot the synthesizer of array 2. The resonances could then be recovered but the wind was too high to start the observation again. We then had another issue with the turbo-pump which went off --> the temperature in the mixing chamber went up to 275 mK. We started again the pump at 20% and the temperature went down too its nominal value. Juan removed the plastic protection surrounding the turbo-pump (maybe the ventilation was not done properly) and we set the pump to 30% again. There are problems with the telescope as well which is blocked even if the wind went down.
At the shift transition period, both telescope and cryostat issues were being resolved; and both were resolved relatively simultaneously. The array cooled back to 205 mK, and resonances were re-acquired. The telescope issue was primarily with the brakes again, but we also encountered an issue at the start of the first tracking scan, but that was resolved. At some point the operator suggested we keep the telescope moving; this was not an issue as we finished the series of X defocus beammaps. After that, we went to HLTau to do some power injection tests; during this period, we kept the telescope moving with many track scans (in between square maps). At the end of this, the operator indicated that things seem to be operating smoothly, and that we need not be as paranoid about constantly moving the telescope (still, we tended to keep the telescope moving). We went back to 3c84 to do a serios of Y defocus maps, ending with 2 nominal focus beammaps. We then went to HATLENS07 for a faint-source (calibration) test, and finished with only 1 large map on Jupiter (not enough time for the second map). We also did a pointing and small map on Ganymede.
While regaining the resonances (around 1:30 UT), we noticed some oscillations at the end of subscans (in the resonances, on the data acquisition software). We did not track this issue very well over the night, but initial impressions suggest that it is (1) always there, and (2) more significant when subscans are done in azimuth (constant elevation).
** NOTE: we REALLY should be skeptical of X & Y focus greater than 0.5 (maybe even less) away from 0. Despite the rather convincing focus plots (https://www.iram.es/mrt-lx1/mrt/ncs/monitor/Nika2Logbook/Scans/20161030s236/plot_20161030s236.png), offline processing reveals that the focus taken (X is in greatest question) is not the right one. ** Suggestion (C. Romero): Limit / strongly favor sources between 40 and 60 degrees in elevation.
Check list before starting the session in the afternoon:
- Cryostat level and tilt to nominal values 9 and 0.5. Did a recover level in the operator consol
- Cryostat temperature of 205 mK as regulated, turbo pump to 30 %
- nika2-a hard disk /home 92 % full (6.2 Tb out of 7.1 Tb), /ssd 16 % full (273 Gb out of 1.8 Tb)
- rsync to nika2-b working
- IMBFITS production started
- Acquisition on: elvin and antenna online, AMC boards working, set monitor autotune
- Cryostat window trap closed, KIDs tuning OK
- Tapas not working (can't save the comments)
Juan disconnected the Mac35 from the NIKA2 network in order to connect it to a monitor next to the observers. Therefore we can make a vnc to the Mac12 from it and see monitor next to the telescope information.
November 1, (Tuesday)
As we observe until later in the morning today, the transition between shifts happened later (towards 2:00 UT). At the start of the second shift, we see that Z focus has been checked recently, and double check the previous X and Y focus measurements. We feel comfortable with the focus as set, and continue to 3c84 to do beammaps (trying to get it at high elevations). We start at nominal focus, and do a series of X-defocus beammaps, and then a series of Y-defocus beammaps. We then plan to do a small pointing session and end with some more large maps on Jupiter. Bye bye NIKA.